Three sides of a triangle are represented by x,x+3 and x+5. The perimeter
of the triangle is 35 units. Solve for x

Answer by checkley77(12844)

x+x+3+x+5=35
3x=35-8
3x=27
x=27/3
x=9 for the short side.
9+3=12 for the middle side.
9+5=14 for the longest side.
Proof:
9+12+14=35
35=35